kevin has 30 nickles and dimes worth $2.40. how many nickles does he have?
Answer by mananth(16949)

nickels = x
dimes = 30-x
..
Value =240 cents
5x+10(30-x)=240
5x+300-10x=240
transpose terms and numbers
10x-5x=300-240
5x=60
/5
x= 12 nickels
dimes = 30-12=18